OpenStudy (anonymous):

The position function of a particle is given by r(t)=⟨4t^2,−4t,t^2+1t⟩. At what time is the speed minimum?

OpenStudy (anonymous):

i have v(t) = <8t, -4, 2t+1>

OpenStudy (anonymous):

no idea what to do now

OpenStudy (anonymous):

This might help. You are given the position vector r(t) = <t^2, 5t, t^2-16t>. Note that r(0)=<0,0,0>, which means the particle starts out at the origin. The velocity vector is v(t)=<2t, 5, 2t-16>. Note that v(0)=<0, 5, -16>, so in fact its initial velocity is nonzero. The speed of the particle is s(t) = sqrt( 4t^2 + 25 + (2t-16)^2 ) = sqrt( 8t^2 -64t + 281) One can check that the speed is always nonzero (either using the formula for s(t), or more simply by examination of v(t) and noting that v(t) is never the zero vector). Minimizing the speed involves finding the critical point for s(t) = sqrt( 8t^2 -64t + 281), which as you note occurs when t=4, although I get s(4) = sqrt(153), not 156.

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

I have s(t) = sqrt(68t^2 + 4t + 17)

OpenStudy (anonymous):

what do I do now? @TruWubs

OpenStudy (anonymous):

how did you find that the critical point is when t = 4?

OpenStudy (perl):

you can take the derivative of the speed s(t)

OpenStudy (anonymous):

then set to 0 and solve for t?

OpenStudy (anonymous):

@perl

OpenStudy (perl):

yes

OpenStudy (anonymous):

ah, thank you :)

OpenStudy (anonymous):

Sorry for not replying Open study is being glitchy he explained fine :)

OpenStudy (anonymous):

np, ty both, yeah I love this site but lots of connection issues with it

OpenStudy (anonymous):

@TruWubs I'm getting s(t) = sqrt(68t^2 + 4t + 17), when i take derivative of that set it to 0 and solve for t i get -1/34, which is incorrect, ive also tried entering 1/34

OpenStudy (anonymous):

I'm sorry I tried again and -1/34 is correct, i was typing it wrong, but how is it possible that its negative time?

OpenStudy (perl):

ok let me check

OpenStudy (anonymous):

i know for sure the answer -1/34 is correct, at least according to the auto grading system, but that makes no sense to me

OpenStudy (anonymous):

okay well

OpenStudy (perl):

Is this r(t) r(t)=<4t^2,−4t,t^2+1t>

OpenStudy (anonymous):

@perl yes that is correct

OpenStudy (anonymous):

lets think about it if u take r'(t), this is giving u the velocity vectors then |r'(t)| is telling u the speed over time

OpenStudy (perl):

I get -1/34

OpenStudy (perl):

so did you take the derivative of s(t) ?

OpenStudy (anonymous):

@perl yes I got -1/34 and its correct, im wondering why a negative time is correct

OpenStudy (anonymous):